OpenStudy (anonymous):

x²+3x+2 divided by x+2 NEED TO SEE WORK PLEASE

OpenStudy (anonymous):

x²+3x+2 divided by x+2 =x+1

OpenStudy (whpalmer4):

To do polynomial long division of the dividend (here \(x^2+3x+2\)) by a divisor (here \(x+2\)), first divide the highest exponent term of the dividend by the highest exponent term of the divisor. \(x^2/x = x\), so the first term of our quotient is \(x\). Now we multiply that first term of the quotient by the entire divisor, write that below our problem, and subtract: x --------------- x+2 | x^2 + 3x + 2 - (x^2 + 2x) ------------- 0x^2 + x + 2 Now we divide the first term of the remaining portion by the first term of the divisor and repeat the process. \(x/x = 1\) so the next term of the quotient is \(1\): x + 1 --------------- x+2 | x^2 + 3x + 2 - (x^2 + 2x) ------------- 0x^2 + x + 2 - ( x + 2) ---------- 0 the remaining portion is 0, so we have no remainder, and our quotient is \(x+1\). Check the result by multiplying the quotient by the divisor and verifying that we get the dividend back: \[(x+2)(x+1) = x(x+1) + 2(x+1) = x^2 + x + 2x + 2 \]\[\qquad= x^2 + 3x + 2\checkmark\]